Fermi National Laboratory
 

Accelerator Update for June 25, 2007 - June 27, 2007

 

Monday June 25

The day shift began with Operations monitoring the Recycle stash (318.9E10), with the Antiproton Source (Pbar) Transport enclosure in access, with D-Zero (D0) and the Collider Detector at Fermilab (CDF) in access.

Other machines are also in controlled access for necessary repairs.

The Pbar rings and Transport accesses were completed by 3:46 PM.

Beam was established to MiniBooNE, NuMI and PBar was stacking by 4:57 PM.

D0 access was completed at 9:06 PM.

Tuesday June 26

The midnight shift began with Operations monitoring the Recycler stash (331.8E10), with the Antiproton Source (Pbar) stacking (67.8E10), with the Linac, Booster, and Main Injector (MI) operating normally, with NuMI and MiniBooNE taking beam, and with the Tev preparing for shot setup.

Operations put the Tevatron (TeV) into shot setup at 12:01 AM.

Operations established store 5516 at 2:13 AM with an initial luminosity of 196.27E30.

Operations reported at 8:15 AM that the Recycler had stashed antiprotons three times during the midnight shift.

At 8:58 AM the MI power supply permit dropped out, experts are investigating.

The MI50 upper power supply was removed from the circuit, Main Injector beam returned at 9:45 AM.

Operations reported at 4:16 PM that the Recycler had stashed antiprotons four times during the day shift.

Operations terminated store 5516 at 9:00 PM.

Operations put the TeV into shot setup at 9:55 PM.

At 11:03 PM the Tevatron quenched during the low beta squeeze, experts are investigating.

Wednesday June 27

The midnight shift began with experts investigating the TeV quench, with the stash at 166.4E10, with Pbar stacking (24.7E10), and with NuMI and MiniBooNE taking beam.

At 2:08 AM Operations got a TeV water leak alarm.

At 3:30 AM experts are repairing the TeV water leak that was traced to the PBar Pre-Vault enclosure. No beam to Pbar during the repair.

PBar returned to stacking at 5:40 AM.

Operations reported at 8:15 AM that the Recycler had stashed twice during the midnight shift.

The Plan for Wednesday

Complete the TeV quench investigation, then put in a Tev store and return to stacking and stashing.


Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 06/27/2007    email Fermilab